OpenStudy (18jonea):

In a survey of 368 people, 81 drink milk straight from the milk carton. For the sample, find the sample proportion, the margin of error, and the interval likely to contain the true population proportion. a. 22%; b. ±11%; c. 11% to 33% a. 28%; b. ±5%; c. 23% to 33% a. 22%; b. ±5%; c. 17% to 27% a. 28%; b. ±11%; c. 17% to 39% is this c or a??

OpenStudy (18jonea):

@jim_thompson5910

OpenStudy (18jonea):

@calculusxy

OpenStudy (18jonea):

@agent0smith

OpenStudy (18jonea):

?

OpenStudy (18jonea):

@agent0smith

OpenStudy (agent0smith):

Sample proportion should be easy. Find the formulas for the ME from your notes or something. Confidence interval formulas.

OpenStudy (18jonea):

i got this 22%; b. ±11%; c. 11% to 33%

OpenStudy (agent0smith):

Show work... i'm not about to do the work to check those answers.

OpenStudy (18jonea):

phat = 81/368 = 0.2201086957 ≈ 0.22 q = 1 - 0.2201086957 = 0.7798913043 = 0.78 P(0.22 - 1.96*sqrt(0.22*0.78/368) < π < 0.22 + 1.96*sqrt(0.22*0.78/368) ) = 0.95 P(0.22 - 0.04 < π < 0.22 + 0.04) = 0.95 P(0.18 < π < 0.26) = 0.95 sample proportion = 0.18 margin of error = 0.04

OpenStudy (18jonea):

so i am guessing on the answer

OpenStudy (agent0smith):

Looks fine.

OpenStudy (18jonea):

so would my answer be correct?

OpenStudy (18jonea):

@agent0smith

OpenStudy (agent0smith):

"Looks fine."

OpenStudy (18jonea):

ok i have 4 more could you help me with them?

OpenStudy (agent0smith):

Possibly. But if you ask "so would I be correct" after I have already implied that you are, I will not help you again. EVER.

OpenStudy (18jonea):

ok